# A Variational Principle for Extended Irreversible Thermodynamics: Heat Conducting Viscous Fluids

### v1

@@ Abstract
Extended irreversible thermodynamics is a theory that extends the classical framework of nonequilibrium thermodynamics by going beyond the local-equilibrium assumption. A notable example of this is the Maxwell-Cattaneo heat flux model, which introduces a time lag in the heat flux response to temperature gradients. In this paper, we develop a variational formulation of the equations of extended irreversible thermodynamics by introducing an action principle for a nonequilibrium Lagrangian that treats thermodynamic fluxes as independent variables. A key feature of this approach is that it naturally extends both Hamilton’s principle of reversible continuum mechanics and the earlier variational formulation of classical irreversible thermodynamics. The variational principle is initially formulated in the material (Lagrangian) description, from which the Eulerian form is derived using material covariance (or relabeling symmetries). The tensorial structure of the thermodynamic fluxes dictates the choice of objective rate in the Eulerian description, leading to the introduction of nonequilibrium stresses arising from both viscous and thermal effects to ensure thermodynamic consistency. This framework naturally results in the Cattaneo-Christov model for heat flux. We also investigate the extension of the approach to accommodate higher-order fluxes and the general form of entropy fluxes. The variational framework presented in this paper has promising applications in the development of structure-preserving and thermodynamically consistent numerical methods. It is particularly relevant for modeling systems where entropy production is a delicate issue that requires careful treatment to ensure consistency with the laws of thermodynamics.
Keywords: Extended irreversible thermodynamics, Heat conducting viscous fluid, Nonequilibrium Lagrangian, Variational formulation, Cattaneo-Christov law.
@@ 1 Introduction
To eliminate these anomalies, a damped version of Fourier’s law, the Maxwell-Cattaneo law, has been proposed by introducing a heat-flux relaxation term, [6, 7, 59]. While the resulting hyperbolic partial differential equation for the temperature solves the above paradox, it turns out that the Maxwell-Cattaneo law is not compatible with the local-equilibrium hypothesis since the classical entropy production computed under this hypothesis fails to be positive. A thermodynamically consistent theory that embodies the Maxwell-Cattaneo law can, however, be constructed that goes beyond the local-equilibrium hypothesis. The main feature of this theory, called extended irreversible thermodynamics [36], is to promote the heat flux, and other thermodynamic fluxes, to the status of independent variables at the same level as other classical variables such as temperature. The first approach to extended irreversible thermodynamics was made by [44, 45] in a classical framework and independently by [34] in a relativistic context. The extension of nonequilibrium thermodynamics beyond the local-equilibrium hypothesis is an extremely lively topic, still under development, with multiple perspectives and a wide range of applications. For an overview, see [35] and the references cited therein. Another prominent approach that extends nonequilibrium thermodynamics beyond the local-equilibrium assumption is rational extended thermodynamics, as discussed in [46, 54]. This approach is based on the closure of moment equations within kinetic theory. We refer to [47], [9], [35], and the references therein for comparisons and discussions of various approaches to nonequilibrium theories
@@
Several variational approaches have been proposed in relation with thermodynamics. At the heart of most of these is the principle of least dissipation of energy, as introduced in [49] and later extended in [50] and [51], which underlies the reciprocal relations in the linear case. Another principle was formulated by [53], [29] as a condition on steady state processes, known as the principle of minimum entropy production. Onsager’s approach was generalized in [61] to the case of systems with nonlinear phenomenological laws. We refer to [31] for reviews and developments of Onsager’s variational principles, as well as for a study of the relationship between Onsager’s and Prigogine’s principles. In this direction, we also refer to e.g. [40, §6] and [33] for overviews on variational approaches to irreversible processes. Unlike the variational approach we develop here, these principles do not seek to extend the Hamilton principle of classical and continuum mechanics. Rather than capturing the complete evolution of a system, they primarily focus on the entropy production equation, aiming to derive the phenomenological laws governing the irreversible processes involved. We refer to [58] and the references therein for a discussion on the construction of variational principles underlying the differential equations of physical theories.
Plan of the paper.
@@
In the remaining part of the introduction we review the variational formulation for finite dimensional thermodynamic systems experiencing mechanical friction and heat conduction, as an extension of Hamilton’s principle and as a preparation for its continuum version revisited in Section 2 for heat conducting viscous fluids in the material description. The variational formulation of classical irreversible thermodynamics in the spatial description is then deduced under the assumption of material covariance. We also recall the general form of the energy and entropy equations for classical irreversible thermodynamics in both the material and spatial description in order to be able to later show the analogies and differences with the case of extended irreversible thermodynamics, see also Table 1. Comments are also given regarding the equivalent use of the first and second Piola-Kirchhoff stress tensors in classical irreversible thermodynamics. In Section 3 we present the variational formulation for extended irreversible thermodynamics in the material description, for continua involving the processes of viscosity and heat conduction. This formulation involves a given nonequilibrium Lagrangian density which depends on the heat and viscous fluxes. The variational formulation is an extension of the one for classical irreversible thermodynamics reviewed in Section 2. The variational formulation in the spatial description is then derived by assuming the material covariance of the nonequilibrium Lagrangian density. The transition from the material to the spatial description directly justifies the choice of one particular objective rate for the heat flux and viscous flux. We also systematically derive from it the momentum equation for extended irreversible thermodynamics in the spatial description, which involves additional nonequilibrium stress tensors. The general form of the energy and entropy equation in the material and spatial description are derived and compared with the corresponding expressions reviewed earlier in the classical irreversible case.
Illustrations.
@@
Figure 1.1 shows the transition from Hamilton’s principle to the proposed variational framework for extended irreversible thermodynamics, passing through the intermediate case of classical irreversible thermodynamics. Here $\mathfrak{L}$ and $\widehat{\mathfrak{L}}$ denote the equilibrium and nonequilibrium Lagrangian densities, respectively. The quantities $J^{\alpha}$ and $X_{\alpha}=\dot{\Lambda}_{\alpha}$ are the thermodynamic fluxes and forces, with $\Lambda_{\alpha}$ being the thermodynamic displacements, associated with with certain irreversible processes $\alpha$. Additionally, $q_{\beta}$ denotes the thermodynamic fluxes on which the nonequilibrium Lagrangian depends. Finally $S$ and $\Sigma$ are the total and internal entropy densities. All notations will be further detailed later. It is evident that the second and third variational formulations are of the d’Alembert type. Table 1 summarizes the energy and entropy balances in each case in the material description, before any phenomenological expressions are chosen.
@@
Lagrangian/HamiltonianReversible ${\begin{array}[]{c}\vspace{0.1cm}\textsf{Lagrangian/Hamiltonian}\\ \vspace{0.1cm}\textsf{Reversible Dynamics}\\ \vspace{0.1cm}\displaystyle\delta\int_{0}^{T}\mathfrak{L}{\rm dt}=0\\ \vspace{0.1cm}\text{No constraints}\end{array}}$lassical ${\begin{array}[]{c}\vspace{0.1cm}\textsf{Classical Irreversible}\\ \vspace{0.1cm}\textsf{Thermodynamics}\\ \vspace{0.1cm}\displaystyle\delta\int_{0}^{T}\mathfrak{L}+(S-\Sigma)\dot{\Gamma}{\rm dt}=0\\ \vspace{0.1cm}\displaystyle\frac{\partial\mathfrak{L}}{\partial S}\dot{\Sigma}=\text{$J^{\alpha}\dot{\Lambda}_{\alpha}$}\\ \vspace{0.1cm}\displaystyle\frac{\partial\mathfrak{L}}{\partial S}\delta\Sigma=\text{$J^{\alpha}\delta\Lambda_{\alpha}$}\end{array}}$xtended ${\begin{array}[]{c}\vspace{0.1cm}\textsf{Extended Irreversible}\\ \vspace{0.1cm}\textsf{Thermodynamics}\\ \vspace{0.1cm}\displaystyle\delta\int_{0}^{T}\widehat{\mathfrak{L}}+(S-\Sigma)\dot{\Gamma}{\rm dt}=0\\ \vspace{0.1cm}\displaystyle\frac{\partial\widehat{\mathfrak{L}}}{\partial S}\dot{\Sigma}+\frac{\partial\widehat{\mathfrak{L}}}{\partial q^{\beta}}\dot{q}^{\beta}=\text{$J^{\alpha}\dot{\Lambda}_{\alpha}$}\\ \vspace{0.1cm}\displaystyle\frac{\partial\widehat{\mathfrak{L}}}{\partial S}\delta\Sigma+\frac{\partial\widehat{\mathfrak{L}}}{\partial q^{\beta}}\delta q^{\beta}=\text{$J^{\alpha}\delta\Lambda_{\alpha}$}\end{array}}$
@@
Table 1: Comparison of energy and entropy balances in reversible dynamics, classical, and extended thermodynamics
The Hamilton principle and its extension to nonequilibrium thermodynamics.
One of the most fundamental statements in classical mechanics is the principle of critical action or Hamilton’s principle, according to which the motion of a mechanical system between two given positions is given by a curve that makes the integral of the Lagrangian of the system critical (see, for instance, [39]):
@@
$L(q,v)=K(q,v)-U(q,S),$
where the kinetic energy $K$ is assumed to be independent of $S$ and $U(q,S)$ is the internal energy, then $T=-\frac{\partial L}{\partial S}=\frac{\partial U}{\partial S}$ recovers the standard definition of the temperature in thermodynamics. We refer to [23, 26] for comments about the structure of the variational principle (1.3)–(1.5), the relation with the Lagrange-d’Alembert principle in nonholonomic mechanics, as well as several applications.
We now write the extension of (1.3)–(1.5) to systems with friction and heat conduction which plays an important guiding role for the present paper, see [23, 26, 28]. Assume that the thermodynamic system is made of several subsystems indexed by $A=1,...,N$, each of them characterized by a configuration variable $q_{A}$ and an entropy variable $S_{A}$. The Lagrangian is hence a function of the form $L(q_{1},...,q_{N},\dot{q}_{1},...,\dot{q}_{N},S_{1},...,S_{N})$. The system involves the following irreversible processes: friction forces $F^{A}$ experienced by subsystem $A$ and heat exchanges between subsystems $A$ and $B$ described by fluxes $J_{AB}$ such that $J_{AB}=J_{BA}$.
To incorporate heat exchange, the new variables $\Gamma^{A}$, $A=1,...,N$ are introduced, called the thermal displacements, such that their time rate of changes coincide with the temperature of each subsystem, $\dot{\Gamma}^{A}=T^{A}$. This is an example of a thermodynamic displacement for an irreversible process, defined so that its time rate of change coincides with thermodynamic force of the process. The introduction of $\Gamma^{A}$ is accompanied with the introduction of an entropy variable $\Sigma_{A}$ whose meaning and importance will be clarified later.
@@
with $\delta q_{A}(0)=\delta q_{A}(T)=0$ and $\delta\Gamma^{A}(0)=\delta\Gamma^{A}(T)=0$, $A=1,...,N$.
@@
for all $A=1,...,N$. In particular, the variations $\delta S_{A}$ and $\delta\Gamma^{A}$ give
@@
Recalling that $-\frac{\partial L}{\partial S_{A}}=T^{A}$ is the temperature of the $A$-th subsystem, its entropy balance is found from (1.10) as
$\dot{S}_{A}=-\frac{1}{T^{A}}F^{A}\cdot\dot{q}_{A}+\sum_{B}J_{AB}\frac{T^{A}-T^{B}}{T^{A}},\quad A=1,...,N.$
@@
$\dot{\Sigma}_{A}=-\frac{1}{T^{A}}F^{A}\cdot\dot{q}_{A}-\sum_{B}J_{AB}\frac{T^{B}}{T^{A}},$
@@ 2.1 Review of the reversible case
Material description.
In the material description the configuration of a continuum, fluid or solid, at time $t\in[0,T]$ is described by a time dependent map $\boldsymbol{\varphi}(t):\mathcal{B}\rightarrow\mathbb{R}^{3}$ assumed in this paper to be a smooth embedding. This map gives the current location $\mathbf{x}=\boldsymbol{\varphi}(t,\mathbf{X})\in\mathbb{R}^{3}$ of each point of the continuum, with $\mathbf{X}\in\mathcal{B}$ the labels of such points. From the configuration map, the Lagrangian velocity $\boldsymbol{V}(t,\mathbf{X})$ and the deformation gradient $\boldsymbol{F}(t,\mathbf{X})$ are defined as
$\boldsymbol{V}(t,\mathbf{X})=\dot{\boldsymbol{\varphi}}(t,\mathbf{X})=\frac{d}{dt}\boldsymbol{\varphi}(t,\mathbf{X}),\qquad\boldsymbol{F}(t,\mathbf{X})=\nabla\boldsymbol{\varphi}(t,\mathbf{X}).$
@@
Variational formulation in the material description.
The equations of evolution follow from Hamilton’s principle
@@
$\varrho\,\ddot{\boldsymbol{\varphi}}=\operatorname{DIV}\big{(}\boldsymbol{P}^{\rm cons}\big{)}\quad\text{on}\quad\mathcal{B},\qquad(\boldsymbol{P}^{\rm cons})^{A}_{a}\boldsymbol{N}_{A}\delta\boldsymbol{\varphi}^{a}=0\quad\text{on}\quad\partial\mathcal{B},$
@@
The notation “cons” refers to a conservative stress tensor to distinguish it from dissipative (irreversible) stress tensors that will be introduced later. For fluids, $\boldsymbol{P}^{\rm cons}$ is related to the pressure, see (2.19) below.
@@
Spatial description.
Besides material frame indifference, isotropic continua also satisfy material covariance, i.e. covariance with respect to diffeomorphisms of $\mathcal{B}$. In terms of the material Lagrangian density $\mathfrak{L}$, material covariance reads
@@
$\mathscr{E}(\nabla(\boldsymbol{\varphi}\circ\boldsymbol{\psi}),(\varrho\circ\boldsymbol{\psi})J_{\boldsymbol{\psi}},(S\circ\boldsymbol{\psi})J_{\boldsymbol{\psi}})=\mathscr{E}(\nabla{\boldsymbol{\varphi}},\varrho,S)\circ\boldsymbol{\psi}$
@@
$\mathscr{E}(\nabla{\boldsymbol{\varphi}},\varrho,S)\varrho=\big{(}\varepsilon(\rho,s)\circ\boldsymbol{\varphi}\big{)}J_{\boldsymbol{\varphi}}$
@@
Variational formulation in the spatial description.
In terms of the spatial variables, Hamilton’s principle (2.3) reads
@@
$\partial_{t}\rho+\operatorname{div}(\rho\boldsymbol{u})=0,\qquad\partial_{t}s+\operatorname{div}(s\boldsymbol{u})=0.$
@@
£$\pounds_{\boldsymbol{u}}\frac{\partial\ell}{\partial\boldsymbol{u}}=\boldsymbol{u}\cdot\nabla\frac{\partial\ell}{\partial\boldsymbol{u}}+\nabla\boldsymbol{u}^{\mathsf{T}}\frac{\partial\ell}{\partial\boldsymbol{u}}+\frac{\partial\ell}{\partial\boldsymbol{u}}\operatorname{div}(\boldsymbol{u}).$ (2.18)
@@
$\rho(\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})=-\nabla p$

### v2

@@ Abstract
Extended irreversible thermodynamics is a theory that expands the classical framework of nonequilibrium thermodynamics by going beyond the local-equilibrium assumption. A notable example of this is the Maxwell-Cattaneo heat flux model, which introduces a time lag in the heat flux response to temperature gradients. In this paper, we develop a variational formulation of the equations of extended irreversible thermodynamics by introducing an action principle for a nonequilibrium Lagrangian that treats thermodynamic fluxes as independent variables. A key feature of this approach is that it naturally extends both Hamilton’s principle of reversible continuum mechanics and the earlier variational formulation of classical irreversible thermodynamics. The variational principle is initially formulated in the material (Lagrangian) description, from which the Eulerian form is derived using material covariance (or relabeling symmetries). The tensorial structure of the thermodynamic fluxes dictates the choice of objective rate in the Eulerian description, and plays a central role in the emergence of nonequilibrium stresses - arising from both viscous and thermal effects - that are essential to ensure thermodynamic consistency. This framework naturally results in the Cattaneo-Christov model for heat flux. We also investigate the extension of the approach to accommodate higher-order fluxes and the general form of entropy fluxes. The variational framework presented in this paper has promising applications in the development of structure-preserving and thermodynamically consistent numerical methods. It is particularly relevant for modeling systems where entropy production is a delicate issue that requires careful treatment to ensure consistency with the laws of thermodynamics.
Keywords: Extended irreversible thermodynamics; Heat conducting viscous fluid; Nonequilibrium Lagrangian; Variational formulation; Cattaneo-Christov law.
@@ 1 Introduction
To eliminate these anomalies, a damped version of Fourier’s law, the Maxwell-Cattaneo law, has been proposed by introducing a heat-flux relaxation term, [6, 7, 67]. While the resulting hyperbolic partial differential equation for the temperature solves the above paradox, it turns out that the Maxwell-Cattaneo law is not compatible with the local-equilibrium hypothesis since the classical entropy production computed under this hypothesis fails to be positive. A thermodynamically consistent theory that embodies the Maxwell-Cattaneo law can, however, be constructed that goes beyond the local-equilibrium hypothesis. The main feature of this theory, called extended irreversible thermodynamics [40], is to promote the heat flux, and other thermodynamic fluxes, to the status of independent variables at the same level as other classical variables such as temperature. The first approach to extended irreversible thermodynamics was made by [48, 49] in a classical framework and independently by [38] in a relativistic context. The extension of nonequilibrium thermodynamics beyond the local-equilibrium hypothesis is an extremely lively topic, still under development, with multiple perspectives and a wide range of applications. For an overview, see [39] and the references cited therein. Another prominent approach that extends nonequilibrium thermodynamics beyond the local-equilibrium assumption is rational extended thermodynamics, as discussed in [50, 59]. This approach is based on the closure of moment equations within kinetic theory. We refer to [51], [9], [39], and the references therein for comparisons and discussions of various approaches to nonequilibrium theories.
@@
Variational principles related to thermodynamics. Several variational approaches have been proposed in relation with thermodynamics. At the heart of most of these is the principle of least dissipation of energy, as introduced in [53] and later extended in [54] and [55], which underlies the reciprocal relations in the linear case. Another principle was formulated by [58], [31] as a condition on steady state processes, known as the principle of minimum entropy production. Onsager’s approach was generalized in [69] to the case of systems with nonlinear phenomenological laws. We refer to [34] for reviews and developments of Onsager’s variational principles, as well as for a study of the relationship between Onsager’s and Prigogine’s principles. In this direction, we also refer to e.g. [44, §6] and [37] for overviews on variational approaches to irreversible processes. Unlike the variational approach we develop here, these principles do not seek to extend the Hamilton principle of classical and continuum mechanics. Rather than capturing the complete evolution of a system, they primarily focus on the entropy production equation, aiming to derive the phenomenological laws governing the irreversible processes involved. We refer to [64] and the references therein for a discussion on the construction of variational principles underlying the differential equations of physical theories.
@@
Plan of the paper. In the remaining part of the introduction we review the variational formulation for finite dimensional thermodynamic systems experiencing mechanical friction and heat conduction, as an extension of Hamilton’s principle and as a preparation for its continuum version revisited in Section 2 for heat conducting viscous fluids in the material description. The variational formulation of classical irreversible thermodynamics in the spatial description is then deduced under the assumption of material covariance. We also recall the general form of the energy and entropy equations for classical irreversible thermodynamics in both the material and spatial description in order to be able to later show the analogies and differences with the case of extended irreversible thermodynamics, see also Table 1. Comments are also given regarding the equivalent use of the first and second Piola-Kirchhoff stress tensors in classical irreversible thermodynamics. In Section 3 we present the variational formulation for extended irreversible thermodynamics in the material description, for continua involving the processes of viscosity and heat conduction. This formulation involves a given nonequilibrium Lagrangian density which depends on the heat and viscous fluxes. The variational formulation is an extension of the one for classical irreversible thermodynamics reviewed in Section 2. The variational formulation in the spatial description is then derived by assuming the material covariance of the nonequilibrium Lagrangian density. The transition from the material to the spatial description directly justifies the choice of one particular objective rate for the heat flux and viscous flux. We also systematically derive from it the momentum equation for extended irreversible thermodynamics in the spatial description, which involves additional nonequilibrium stress tensors. The general form of the energy and entropy equation in the material and spatial description are derived and compared with the corresponding expressions reviewed earlier in the classical irreversible case.
@@
Illustrations. Figure 1.1 illustrates the transition from Hamilton’s principle to the proposed variational framework for extended irreversible thermodynamics, passing through the intermediate case of classical irreversible thermodynamics. Here $\mathfrak{L}$ and $\widehat{\mathfrak{L}}$ denote the equilibrium and nonequilibrium Lagrangian densities, respectively. The quantities $J^{\alpha}$ and $X_{\alpha}=\dot{\Lambda}_{\alpha}$ are the thermodynamic fluxes and forces, with $\Lambda_{\alpha}$ being the thermodynamic displacements, associated with with certain irreversible processes $\alpha$. Additionally, $q_{\beta}$ denotes the thermodynamic fluxes on which the nonequilibrium Lagrangian depends. Finally $S$ and $\Sigma$ are the total and internal entropy densities. All notations will be further detailed later. It is evident that the second and third variational formulations are of the d’Alembert type. Table 1 summarizes the energy and entropy balances in each case in the material description, before any phenomenological expressions are chosen.
@@
${\begin{array}[]{c}\vskip 2.84544pt\textsf{Lagrangian/Hamiltonian}\\ \vskip 2.84544pt\textsf{Reversible Dynamics}\\ \vskip 2.84544pt\displaystyle\delta\int_{0}^{T}\mathfrak{L}{\rm dt}=0\\ \vskip 2.84544pt\text{No constraints}\end{array}}$${\begin{array}[]{c}\vskip 2.84544pt\textsf{Classical Irreversible}\\ \vskip 2.84544pt\textsf{Thermodynamics}\\ \vskip 2.84544pt\displaystyle\delta\int_{0}^{T}\mathfrak{L}+(S-\Sigma)\dot{\Gamma}{\rm dt}=0\\ \vskip 2.84544pt\displaystyle\frac{\partial\mathfrak{L}}{\partial S}\dot{\Sigma}=\text{$J^{\alpha}\dot{\Lambda}_{\alpha}$}\\ \vskip 2.84544pt\displaystyle\frac{\partial\mathfrak{L}}{\partial S}\delta\Sigma=\text{$J^{\alpha}\delta\Lambda_{\alpha}$}\end{array}}$${\begin{array}[]{c}\vskip 2.84544pt\textsf{Extended Irreversible}\\ \vskip 2.84544pt\textsf{Thermodynamics}\\ \vskip 2.84544pt\displaystyle\delta\int_{0}^{T}\widehat{\mathfrak{L}}+(S-\Sigma)\dot{\Gamma}{\rm dt}=0\\ \vskip 2.84544pt\displaystyle\frac{\partial\widehat{\mathfrak{L}}}{\partial S}\dot{\Sigma}+\frac{\partial\widehat{\mathfrak{L}}}{\partial q^{\beta}}\dot{q}^{\beta}=\text{$J^{\alpha}\dot{\Lambda}_{\alpha}$}\\ \vskip 2.84544pt\displaystyle\frac{\partial\widehat{\mathfrak{L}}}{\partial S}\delta\Sigma+\frac{\partial\widehat{\mathfrak{L}}}{\partial q^{\beta}}\delta q^{\beta}=\text{$J^{\alpha}\delta\Lambda_{\alpha}$}\end{array}}$
@@
Table 1: Comparison of energy and entropy balances in reversible dynamics, classical, and extended thermodynamics (material frame description).
The Hamilton principle and its extension to nonequilibrium thermodynamics. One of the most fundamental statements in classical mechanics is the principle of critical action or Hamilton’s principle, according to which the motion of a mechanical system between two given positions is given by a curve that makes the integral of the Lagrangian of the system critical (see, for instance, [43]):
@@
L(q,v,S)=K(q,v)U(q,S),L(q,v,S)=K(q,v)-U(q,S), (1.7)
where the kinetic energy $K$ is assumed to be independent of $S$ and $U(q,S)$ is the internal energy, then $T=-\frac{\partial L}{\partial S}=\frac{\partial U}{\partial S}$ recovers the standard definition of the temperature in thermodynamics. We refer to [25, 28] for comments about the structure of the variational principle (1.3)–(1.5), the relation with the Lagrange-d’Alembert principle in nonholonomic mechanics, as well as several applications. We refer the reader to [20, Remark 2.2] for a discussion of the difference between this variational formulation and the variational-like approach based on Rayleigh dissipation functions.
We now write the extension of (1.3)–(1.5) to systems with friction and heat conduction which plays an important guiding role for the present paper, see [25, 28, 30]. Assume that the thermodynamic system is made of several subsystems indexed by A=1,,NA=1,...,N, each of them characterized by a configuration variable $q_{A}$ and an entropy variable $S_{A}$. The Lagrangian is hence a function of the form $L(q_{1},...,q_{N},\dot{q}_{1},...,\dot{q}_{N},S_{1},...,S_{N})$. The system involves the following irreversible processes: friction forces $F^{A}$ experienced by subsystem $A$ and heat exchanges between subsystems $A$ and $B$ described by fluxes $J_{AB}$ such that $J_{AB}=J_{BA}$.
To incorporate heat exchange, the new variables $\Gamma^{A}$, A=1,,NA=1,...,N are introduced, called the thermal displacements, such that their time rate of changes coincide with the temperature of each subsystem, $\dot{\Gamma}^{A}=T^{A}$, see also Remark 2.6. This is an example of a thermodynamic displacement for an irreversible process, introduced in [25, 28], defined so that its time rate of change coincides with thermodynamic force of the process. The introduction of $\Gamma^{A}$ is accompanied with the introduction of an entropy variable $\Sigma_{A}$ whose meaning and importance will be clarified later.
@@
with $\delta q_{A}(0)=\delta q_{A}(T)=0$ and $\delta\Gamma^{A}(0)=\delta\Gamma^{A}(T)=0$, A=1,,NA=1,...,N.
@@
for all A=1,,NA=1,...,N. In particular, the variations $\delta S_{A}$ and $\delta\Gamma^{A}$ give
@@
Recalling that $\frac{\partial L}{\partial S_{A}}=T^{A}$ is the temperature of the $A$-th subsystem, its entropy balance is found from (1.11) as
$\dot{S}_{A}=-\frac{1}{T^{A}}F^{A}\cdot\dot{q}_{A}+\sum_{B}J_{AB}\frac{T^{A}-T^{B}}{T^{A}},\quad A=1,...,N.$ (1.13)
@@
$\dot{\Sigma}_{A}=-\frac{1}{T^{A}}F^{A}\cdot\dot{q}_{A}-\sum_{B}J_{AB}\frac{T^{B}}{T^{A}},$ (1.14)
@@ 2.1 Review of the reversible case
Material description. In the material description the configuration of a continuum, fluid or solid, at time $t\in[0,T]$ is described by a time dependent map $\boldsymbol{\varphi}(t):\mathcal{B}\rightarrow\mathbb{R}^{3}$ assumed in this paper to be a smooth embedding. This map gives the current location $\mathbf{x}=\boldsymbol{\varphi}(t,\mathbf{X})\in\mathbb{R}^{3}$ of each point of the continuum, with $\mathbf{X}\in\mathcal{B}$ the labels of such points. From the configuration map, the Lagrangian velocity $\boldsymbol{V}(t,\mathbf{X})$ and the deformation gradient $\boldsymbol{F}(t,\mathbf{X})$ are defined as
$\boldsymbol{V}(t,\mathbf{X})=\dot{\boldsymbol{\varphi}}(t,\mathbf{X})=\frac{d}{dt}\boldsymbol{\varphi}(t,\mathbf{X}),\qquad\boldsymbol{F}(t,\mathbf{X})=\nabla\boldsymbol{\varphi}(t,\mathbf{X}).$ (2.1)
@@
Variational formulation in the material description. The equations of evolution follow from Hamilton’s principle
@@
$\varrho\,\ddot{\boldsymbol{\varphi}}=\operatorname{DIV}\big{(}\boldsymbol{P}^{\rm cons}\big{)}\quad\text{on}\quad\mathcal{B},\qquad(\boldsymbol{P}^{\rm cons})^{A}_{a}\boldsymbol{N}_{A}\delta\boldsymbol{\varphi}^{a}=0\quad\text{on}\quad\partial\mathcal{B},$ (2.7)
@@
The notation ”cons” refers to a conservative stress tensor to distinguish it from dissipative (irreversible) stress tensors that will be introduced later. For fluids, $\boldsymbol{P}^{\rm cons}$ is related to the pressure, see (2.25) below.
@@
Spatial description. Besides material frame indifference, isotropic continua also satisfy material covariance, i.e. covariance with respect to diffeomorphisms of $\mathcal{B}$. In terms of the material Lagrangian density $\mathfrak{L}$, material covariance reads
@@
$\mathscr{E}(\nabla(\boldsymbol{\varphi}\circ\boldsymbol{\psi}),(\varrho\circ\boldsymbol{\psi})J_{\boldsymbol{\psi}},(S\circ\boldsymbol{\psi})J_{\boldsymbol{\psi}})=\mathscr{E}(\nabla{\boldsymbol{\varphi}},\varrho,S)\circ\boldsymbol{\psi}$ (2.15)
@@
$\mathscr{E}(\nabla{\boldsymbol{\varphi}},\varrho,S)\varrho=\big{(}\varepsilon(\rho,s)\circ\boldsymbol{\varphi}\big{)}J_{\boldsymbol{\varphi}}$ (2.16)
@@
Variational formulation in the spatial description. In terms of the spatial variables, Hamilton’s principle (2.4) reads
@@
$\partial_{t}\rho+\operatorname{div}(\rho\boldsymbol{u})=0,\qquad\partial_{t}s+\operatorname{div}(s\boldsymbol{u})=0.$ (2.22)
@@
$\pounds_{\boldsymbol{u}}\frac{\partial\ell}{\partial\boldsymbol{u}}=\boldsymbol{u}\cdot\nabla\frac{\partial\ell}{\partial\boldsymbol{u}}+\nabla\boldsymbol{u}^{\mathsf{T}}\frac{\partial\ell}{\partial\boldsymbol{u}}+\frac{\partial\ell}{\partial\boldsymbol{u}}\operatorname{div}(\boldsymbol{u}).$ (2.23)
@@
$\rho(\partial_{t}\boldsymbol{u}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})=-\nabla p$ (2.24)